From the same store, Anne bought 3 loaves of bread and 2 bottles of milk for $15.30, and Joe bought 2 loaves of bread and a bottle of milk for $9.20. What is the price for one of each?
Answer by Nate(3500)

-3b - 2m = -15.3
+
4b + 2m = 18.4
=
b = 3.10
plug ....
2b + m = 9.2
2(3.1) + m = 9.2
6.2 + m = 9.2
m = 3.00
The loaves of bread are $3.10 each, but the bottles of milk are $3.00 each.